OpenStudy (anonymous):

sinv=1/4 what is cos(2v)?

OpenStudy (eyust707):

do you know a way to express sine in terms of cosine?

OpenStudy (anonymous):

no haha

OpenStudy (anonymous):

not too sure how to go about this?

OpenStudy (experimentx):

Use these steps \[ \cos x = \sqrt{1-\sin^2x} \] \[ \cos^2 x = 1-\sin^2 x \] \[ \cos 2x = 1 - 2 \sin^2x \]

OpenStudy (anonymous):

cos2x=1-(1/16)

OpenStudy (anonymous):

well... cos2x=1-2sin1/16?

OpenStudy (anonymous):

\[\cos \left(2 \ \sin ^{-1}\left(\frac{1}{4}\right)\right)=0.875 \]

OpenStudy (experimentx):

cos2x=1-(2/16) ^